Find the midpoint of each segment with the given endpoints.
C(–5, –1) and D(–7, 7)

Found 2 solutions by Alan3354, Chris G:
Answer by Alan3354(69443)

Find the average of x and y separately.
For x: (-5 -7)/2 = -6 (you could eyeball that one)
For y: (-1 +7)/2 = 3
Midpoint at (-6,3)

Answer by Chris G(2)